remi2019:

Solve for x: 1/8 (8x + 15) = 24 plz help I am so confused

janiya101:

ok so what do you think for your awnser

janiya101:

1/8 * (8x + 15) = 24 1/8 * 8x + 1/8 * 15 = 24 x + 15/8 = 24 x = 24 - 15/8 x = 192/8 - 15/8 = 177/8 = 22.125 The correct result would be 177/8.

supie:

\(\LARGE\frac18(8x+15)=24\) Simplify; distribute: \(\LARGE(\frac18)(8x)+(\frac18)(15)=24\) \(\LARGE x+\frac{15}{8}=24\) \(\frac {-15}{8}\) both sides: \(\LARGE x+\frac{15}{8}−\frac{15}{8}=24−\frac{15}{8}\) \(\LARGE x= \frac{?}{?}\)

remi2019:

thx

remi2019:

i did thx